# Efficient Model-Stealing Attacks Against Inductive Graph Neural Networks

### v2

@@ 6 Experimental Evaluation
In this section, we conduct a thorough experimental analysis of a model stealing attack on inductive GNNs and show that our approach outperforms the state-of-the-art algorithm by Shen et al. [30]. In what follows, we introduce the experimental setup, evaluate Type I and Type II attacks from theft and reconnaissance perspectives, and investigate the impact of different query budgets on attack performance. To our knowledge, ours and Shen et al. [30] are the only existing model stealing methods applicable to inductive-GNN.
6.1 Experimental Setup
Datasets We evaluate performance of our attack on 6 benchmark datasets for evaluating GNN performance [15, 10, 36]: DBLP [24], Pubmed [27], Citeseer Full (referred to as Citeseer) [8], Coauthor Physics (referred to as Coauthor) [29], ACM [35], and Amazon Co-purchase Network for Photos (referred to as Amazon) [22]. Pubmed, and Citeseer are citation networks, with nodes representing publications and edges denoting citations. We leverage these datasets to assess our attack efficacy across varied graph characteristics such as size, node features, and classes. For each dataset, we divide it into three segments. The initial 20% involves randomly selected nodes for training the target model $\mathcal{M}_{T}$. The subsequent 30% forms our query graph $\mathbf{G}_{Q}$. We demonstrate the continued effectiveness of our attacks even with a reduced number of nodes in the query graph (refer to Section 6.4). The remaining 50% constitutes the testing data for both $\mathcal{M}_{T}$ and $\mathcal{M}_{S}$. Noteworthy, in the official implementation of [30] the testing data is sampled from the whole dataset, creating an overlap between the first two and the third data split. We conduct all experiments, including the state-of-the-art baseline by Shen et al. [30] on disjoint data sets.
@@ 6.1 Experimental Setup
Metrics. Following  [13] taxonomy, we employ two metrics—accuracy and fidelity—to assess our attack’s performance. For the theft adversary, we gauge accuracy (correct predictions divided by total predictions), a widely-used metric in GNN node classification evaluation [15, 10, 33]. As the reconnaissance adversary aims to mimic the target model’s behavior, we utilize fidelity (predictions agreed upon by both $\mathcal{M}_{S}$ and $\mathcal{M}_{T}$) as our second evaluation metric [14, 13].
@@ 6.2 Evaluation of Type I Attacks
Dataset $\mathcal{M}_{T}$
GIN GAT SAGE
DBLP 0.772 0.759 0.781
Pubmed 0.844 0.821 0.856
Citeseer 0.834 0.836 0.838
Coauthor 0.912 0.941 0.943
ACM 0.890 0.897 0.905
Amazon 0.819 0.920 0.902
@@
Accuracy Fidelity Accuracy Fidelity Accuracy Fidelity
Shen et al. 0.738$\pm$0.001 0.805$\pm$0.000 0.715$\pm$0.003 0.782$\pm$0.004 0.735$\pm$0.001 0.799$\pm$0.002
Prediction ours 0.774$\pm$0.002 0.826$\pm$0.002 0.776$\pm$0.001 0.847$\pm$0.002 0.781$\pm$0.001 0.834$\pm$0.001
Shen et al. 0.705$\pm$0.007 0.751$\pm$0.000 0.695$\pm$0.016 0.757$\pm$0.020 0.704$\pm$0.031 0.755$\pm$0.038
Projection ours 0.740$\pm$0.001 0.783$\pm$0.000 0.721$\pm$0.004 0.772$\pm$0.003 0.746$\pm$0.002 0.788$\pm$0.001
Shen et al. 0.705$\pm$0.002 0.759$\pm$0.000 0.683$\pm$0.006 0.723$\pm$0.013 0.710$\pm$0.002 0.762$\pm$0.001
DBLP Embedding ours 0.770$\pm$0.000 0.806$\pm$0.004 0.774$\pm$0.004 0.827$\pm$0.005 0.755$\pm$0.009 0.796$\pm$0.005
Shen et al. 0.846$\pm$0.001 0.912$\pm$0.001 0.825$\pm$0.002 0.906$\pm$0.001 0.846$\pm$0.001 0.914$\pm$0.000
Prediction ours 0.856$\pm$0.001 0.892$\pm$0.001 0.845$\pm$0.001 0.897$\pm$0.001 0.859$\pm$0.002 0.896$\pm$0.002
Shen et al. 0.816$\pm$0.026 0.872$\pm$0.029 0.797$\pm$0.037 0.864$\pm$0.052 0.793$\pm$0.016 0.848$\pm$0.020
Projection ours 0.848$\pm$0.001 0.883$\pm$0.000 0.840$\pm$0.002 0.896$\pm$0.001 0.855$\pm$0.002 0.889$\pm$0.002
Shen et al. 0.853$\pm$0.002 0.910$\pm$0.001 0.834$\pm$0.001 0.902$\pm$0.005 0.846$\pm$0.001 0.911$\pm$0.003
Pubmed Embedding ours 0.863$\pm$0.002 0.888$\pm$0.003 0.853$\pm$0.001 0.898$\pm$0.003 0.862$\pm$0.003 0.879$\pm$0.005
Shen et al. 0.773$\pm$0.001 0.818$\pm$0.001 0.725$\pm$0.007 0.766$\pm$0.009 0.762$\pm$0.002 0.794$\pm$0.003
Prediction ours 0.820$\pm$0.003 0.843$\pm$0.001 0.827$\pm$0.003 0.861$\pm$0.003 0.814$\pm$0.003 0.843$\pm$0.002
Shen et al. 0.668$\pm$0.050 0.688$\pm$0.051 0.704$\pm$0.012 0.733$\pm$0.015 0.664$\pm$0.043 0.695$\pm$0.053
Projection ours 0.751$\pm$0.003 0.763$\pm$0.002 0.766$\pm$0.005 0.779$\pm$0.004 0.631$\pm$0.006 0.648$\pm$0.008
Shen et al. 0.736$\pm$0.005 0.772$\pm$0.004 0.730$\pm$0.004 0.771$\pm$0.007 0.753$\pm$0.003 0.797$\pm$0.002
Citeseer Embedding ours 0.810$\pm$0.003 0.827$\pm$0.001 0.828$\pm$0.000 0.852$\pm$0.000 0.813$\pm$0.004 0.818$\pm$0.011
Shen et al. 0.869$\pm$0.002 0.808$\pm$0.000 0.759$\pm$0.021 0.722$\pm$0.028 0.842$\pm$0.012 0.801$\pm$0.016
Prediction ours 0.915$\pm$0.003 0.839$\pm$0.002 0.855$\pm$0.006 0.825$\pm$0.016 0.881$\pm$0.009 0.816$\pm$0.005
Shen et al. 0.609$\pm$0.055 0.633$\pm$0.100 0.606$\pm$0.052 0.600$\pm$0.081 0.635$\pm$0.021 0.646$\pm$0.011
Projection ours 0.635$\pm$0.002 0.616$\pm$0.007 0.619$\pm$0.041 0.591$\pm$0.042 0.524$\pm$0.015 0.538$\pm$0.014
Shen et al. 0.891$\pm$0.005 0.820$\pm$0.000 0.777$\pm$0.032 0.737$\pm$0.024 0.877$\pm$0.004 0.809$\pm$0.003
Amazon Embedding ours 0.919$\pm$0.000 0.828$\pm$0.005 0.901$\pm$0.009 0.825$\pm$0.010 0.893$\pm$0.007 0.823$\pm$0.002
Shen et al. 0.913$\pm$0.001 0.930$\pm$0.000 0.864$\pm$0.009 0.888$\pm$0.010 0.901$\pm$0.006 0.918$\pm$0.007
Prediction ours 0.942$\pm$0.002 0.950$\pm$0.002 0.921$\pm$0.002 0.942$\pm$0.003 0.942$\pm$0.000 0.957$\pm$0.001
Shen et al. 0.790$\pm$0.047 0.797$\pm$0.000 0.832$\pm$0.018 0.854$\pm$0.018 0.773$\pm$0.068 0.784$\pm$0.070
Projection ours 0.920$\pm$0.004 0.926$\pm$0.004 0.881$\pm$0.013 0.896$\pm$0.012 0.824$\pm$0.008 0.836$\pm$0.008
Shen et al. 0.887$\pm$0.004 0.902$\pm$0.000 0.878$\pm$0.000 0.891$\pm$0.002 0.910$\pm$0.005 0.922$\pm$0.005
Coauthor Embedding ours 0.943$\pm$0.003 0.948$\pm$0.005 0.934$\pm$0.001 0.944$\pm$0.002 0.941$\pm$0.001 0.941$\pm$0.002
Shen et al. 0.733$\pm$0.007 0.759$\pm$0.007 0.725$\pm$0.040 0.744$\pm$0.051 0.750$\pm$0.008 0.793$\pm$0.007
Prediction ours 0.895$\pm$0.005 0.913$\pm$0.006 0.888$\pm$0.008 0.934$\pm$0.008 0.895$\pm$0.004 0.928$\pm$0.004
Shen et al. 0.794$\pm$0.009 0.810$\pm$0.016 0.780$\pm$0.028 0.791$\pm$0.036 0.801$\pm$0.039 0.831$\pm$0.042
Projection ours 0.770$\pm$0.011 0.798$\pm$0.008 0.757$\pm$0.004 0.793$\pm$0.003 0.816$\pm$0.016 0.823$\pm$0.020
Shen et al. 0.811$\pm$0.010 0.836$\pm$0.008 0.817$\pm$0.010 0.840$\pm$0.011 0.835$\pm$0.003 0.865$\pm$0.003
ACM Embedding ours 0.842$\pm$0.007 0.822$\pm$0.015 0.857$\pm$0.009 0.870$\pm$0.009 0.882$\pm$0.009 0.894$\pm$0.011

### v4

@@ 6 Experimental Evaluation
In this section, we conduct a thorough experimental analysis of a model stealing attack on inductive GNNs and show that our approach outperforms the state-of-the-art algorithm by Shen et al. [33]. In what follows, we introduce the experimental setup, evaluate Type I and Type II attacks from theft and reconnaissance perspectives, and investigate the impact of different query budgets on attack performance. To our knowledge, ours and Shen et al. [33] are the only existing model stealing methods applicable to inductive GNN.
Datasets. We evaluate performance of our attack on 6 benchmark datasets for evaluating GNN performance [16, 11, 40]: DBLP [25], Pubmed [30], Citeseer Full (referred to as Citeseer) [9], Coauthor Physics (referred to as Coauthor) [32], ACM [38], and Amazon Co-purchase Network for Photos (referred to as Amazon) [23]. We use these datasets to evaluate our attack’s efficacy across different graph characteristics such as size, node features, and classes. Each dataset is split into three segments: the first 20% of randomly selected nodes for training the target model $\mathcal{M}_{T}$, the next 30% as the query graph $\mathbf{G}_{Q}$, and the final 50% as testing data for both $\mathcal{M}_{T}$ and $\mathcal{M}_{S}$. Further, we show that our attacks remain effective even with a reduced number of nodes in the query graph (see Section 6.3). Importantly, in the official implementation of [33] the testing data is sampled from the whole dataset, creating an overlap between the first two and the third data split. We conduct all experiments, including the state-of-the-art baseline by Shen et al. [33] on disjoint data sets.
@@
Metrics. Following  [14] taxonomy, we employ three metrics — accuracy, F1 score and fidelity — to assess our attack’s performance. For the theft adversary, we gauge accuracy (correct predictions divided by total predictions) and F1 score (the harmonic mean of precision and recall), a widely-used metric in GNN node classification evaluation [16, 11, 36]. As the reconnaissance adversary aims to mimic the target model’s behavior, we utilize fidelity (predictions agreed upon by both $\mathcal{M}_{S}$ and $\mathcal{M}_{T}$) as our evaluation metric [15, 14].
Dataset GIN GAT SAGE
Accuracy F1 Accuracy F1 Accuracy F1
DBLP 0.772 0.713 0.759 0.693 0.781 0.715
Pubmed 0.844 0.839 0.821 0.816 0.856 0.852
Citeseer 0.834 0.836 0.836 0.838 0.838 0.839
Coauthor 0.912 0.907 0.941 0.923 0.943 0.924
ACM 0.890 0.890 0.897 0.896 0.905 0.905
Amazon 0.819 0.754 0.920 0.912 0.902 0.897
@@ 6.1 Evaluation of Type I Attacks
Accuracy Fidelity F1 score Accuracy Fidelity F1 score Accuracy Fidelity F1 score
Shen et al. 0.738$\pm$0.001 0.805$\pm$0.000 0.675$\pm$0.002 0.715$\pm$0.003 0.782$\pm$0.004 0.620$\pm$0.002 0.735$\pm$0.001 0.799$\pm$0.002 0.679$\pm$0.001
Prediction ours 0.774$\pm$0.002 0.826$\pm$0.002 0.718$\pm$0.002 0.776$\pm$0.001 0.847$\pm$0.002 0.717$\pm$0.002 0.781$\pm$0.001 0.834$\pm$0.001 0.723$\pm$0.002
Shen et al. 0.705$\pm$0.007 0.751$\pm$0.000 0.569$\pm$0.004 0.695$\pm$0.016 0.757$\pm$0.020 0.549$\pm$0.011 0.704$\pm$0.031 0.755$\pm$0.038 0.588$\pm$0.022
Projection ours 0.740$\pm$0.001 0.783$\pm$0.000 0.607$\pm$0.004 0.721$\pm$0.004 0.772$\pm$0.003 0.593$\pm$0.001 0.746$\pm$0.002 0.788$\pm$0.001 0.579$\pm$0.004
Shen et al. 0.705$\pm$0.002 0.759$\pm$0.000 0.626$\pm$0.001 0.683$\pm$0.006 0.723$\pm$0.013 0.621$\pm$0.001 0.710$\pm$0.002 0.762$\pm$0.001 0.645$\pm$0.004
DBLP Embedding ours 0.770$\pm$0.000 0.806$\pm$0.004 0.804$\pm$0.001 0.774$\pm$0.004 0.827$\pm$0.005 0.836$\pm$0.002 0.755$\pm$0.009 0.796$\pm$0.005 0.787$\pm$0.002
Shen et al. 0.846$\pm$0.001 0.912$\pm$0.001 0.843$\pm$0.004 0.825$\pm$0.002 0.906$\pm$0.001 0.819$\pm$0.005 0.846$\pm$0.001 0.914$\pm$0.000 0.840$\pm$0.005
Prediction ours 0.856$\pm$0.001 0.892$\pm$0.001 0.847$\pm$0.004 0.845$\pm$0.001 0.897$\pm$0.001 0.835$\pm$0.004 0.859$\pm$0.002 0.896$\pm$0.002 0.852$\pm$0.001
Shen et al. 0.816$\pm$0.026 0.872$\pm$0.029 0.795$\pm$0.005 0.797$\pm$0.037 0.864$\pm$0.052 0.808$\pm$0.005 0.793$\pm$0.016 0.848$\pm$0.020 0.833$\pm$0.005
Projection ours 0.848$\pm$0.001 0.883$\pm$0.000 0.841$\pm$0.009 0.840$\pm$0.002 0.896$\pm$0.001 0.824$\pm$0.004 0.855$\pm$0.002 0.889$\pm$0.002 0.840$\pm$0.002
Shen et al. 0.853$\pm$0.002 0.910$\pm$0.001 0.847$\pm$0.002 0.834$\pm$0.001 0.902$\pm$0.005 0.829$\pm$0.002 0.846$\pm$0.001 0.911$\pm$0.003 0.839$\pm$0.002
Pubmed Embedding ours 0.863$\pm$0.002 0.888$\pm$0.003 0.857$\pm$0.004 0.853$\pm$0.001 0.898$\pm$0.003 0.850$\pm$0.009 0.862$\pm$0.003 0.879$\pm$0.005 0.862$\pm$0.005
Shen et al. 0.773$\pm$0.001 0.818$\pm$0.001 0.777$\pm$0.002 0.725$\pm$0.007 0.766$\pm$0.009 0.741$\pm$0.001 0.762$\pm$0.002 0.794$\pm$0.003 0.768$\pm$0.001
Prediction ours 0.820$\pm$0.003 0.843$\pm$0.001 0.820$\pm$0.003 0.827$\pm$0.003 0.861$\pm$0.003 0.827$\pm$0.005 0.814$\pm$0.003 0.843$\pm$0.002 0.814$\pm$0.009
Shen et al. 0.668$\pm$0.050 0.688$\pm$0.051 0.681$\pm$0.001 0.704$\pm$0.012 0.733$\pm$0.015 0.754$\pm$0.005 0.664$\pm$0.043 0.695$\pm$0.053 0.599$\pm$0.002
Projection ours 0.751$\pm$0.003 0.763$\pm$0.002 0.740$\pm$0.001 0.766$\pm$0.005 0.779$\pm$0.004 0.760$\pm$0.009 0.631$\pm$0.006 0.648$\pm$0.008 0.612$\pm$0.001
Shen et al. 0.736$\pm$0.005 0.772$\pm$0.004 0.748$\pm$0.004 0.730$\pm$0.004 0.771$\pm$0.007 0.782$\pm$0.001 0.753$\pm$0.003 0.797$\pm$0.002 0.745$\pm$0.001
Citeseer Embedding ours 0.810$\pm$0.003 0.827$\pm$0.001 0.815$\pm$0.001 0.828$\pm$0.000 0.852$\pm$0.000 0.837$\pm$0.002 0.813$\pm$0.004 0.818$\pm$0.011 0.810$\pm$0.001
Shen et al. 0.869$\pm$0.002 0.808$\pm$0.000 0.821$\pm$0.001 0.759$\pm$0.021 0.722$\pm$0.028 0.671$\pm$0.010 0.842$\pm$0.012 0.801$\pm$0.016 0.789$\pm$0.002
Prediction ours 0.915$\pm$0.003 0.839$\pm$0.002 0.877$\pm$0.003 0.855$\pm$0.006 0.825$\pm$0.016 0.796$\pm$0.003 0.881$\pm$0.009 0.816$\pm$0.005 0.856$\pm$0.001
Shen et al. 0.609$\pm$0.055 0.633$\pm$0.100 0.512$\pm$0.012 0.606$\pm$0.052 0.600$\pm$0.081 0.559$\pm$0.001 0.635$\pm$0.021 0.646$\pm$0.011 0.540$\pm$0.010
Projection ours 0.635$\pm$0.002 0.616$\pm$0.007 0.673$\pm$0.002 0.619$\pm$0.041 0.591$\pm$0.042 0.564$\pm$0.009 0.524$\pm$0.015 0.538$\pm$0.014 0.613$\pm$0.002
Shen et al. 0.891$\pm$0.005 0.820$\pm$0.000 0.866$\pm$0.004 0.777$\pm$0.032 0.737$\pm$0.024 0.781$\pm$0.005 0.877$\pm$0.004 0.809$\pm$0.003 0.822$\pm$0.001
Amazon Embedding ours 0.919$\pm$0.000 0.828$\pm$0.005 0.887$\pm$0.001 0.901$\pm$0.009 0.825$\pm$0.010 0.834$\pm$0.001 0.893$\pm$0.007 0.823$\pm$0.002 0.861$\pm$0.004
Shen et al. 0.913$\pm$0.001 0.930$\pm$0.000 0.888$\pm$0.004 0.864$\pm$0.009 0.888$\pm$0.010 0.844$\pm$0.001 0.901$\pm$0.006 0.918$\pm$0.007 0.885$\pm$0.002
Prediction ours 0.942$\pm$0.002 0.950$\pm$0.002 0.926$\pm$0.009 0.921$\pm$0.002 0.942$\pm$0.003 0.876$\pm$0.003 0.942$\pm$0.000 0.957$\pm$0.001 0.923$\pm$0.009
Shen et al. 0.790$\pm$0.047 0.797$\pm$0.000 0.760$\pm$0.005 0.832$\pm$0.018 0.854$\pm$0.018 0.725$\pm$0.001 0.773$\pm$0.068 0.784$\pm$0.070 0.655$\pm$0.001
Projection ours 0.920$\pm$0.004 0.926$\pm$0.004 0.888$\pm$0.002 0.881$\pm$0.013 0.896$\pm$0.012 0.866$\pm$0.001 0.824$\pm$0.008 0.836$\pm$0.008 0.675$\pm$0.009
Shen et al. 0.887$\pm$0.004 0.902$\pm$0.000 0.852$\pm$0.004 0.878$\pm$0.000 0.891$\pm$0.002 0.869$\pm$0.002 0.910$\pm$0.005 0.922$\pm$0.005 0.882$\pm$0.004
Coauthor Embedding ours 0.943$\pm$0.003 0.948$\pm$0.005 0.926$\pm$0.001 0.934$\pm$0.001 0.944$\pm$0.002 0.911$\pm$0.004 0.941$\pm$0.001 0.941$\pm$0.002 0.915$\pm$0.005
Shen et al. 0.733$\pm$0.007 0.759$\pm$0.007 0.736$\pm$0.004 0.725$\pm$0.040 0.744$\pm$0.051 0.741$\pm$0.004 0.750$\pm$0.008 0.793$\pm$0.007 0.738$\pm$0.004
Prediction ours 0.895$\pm$0.005 0.913$\pm$0.006 0.896$\pm$0.001 0.888$\pm$0.008 0.934$\pm$0.008 0.889$\pm$0.003 0.895$\pm$0.004 0.928$\pm$0.004 0.906$\pm$0.005
Shen et al. 0.794$\pm$0.009 0.810$\pm$0.016 0.712$\pm$0.004 0.780$\pm$0.028 0.791$\pm$0.036 0.737$\pm$0.005 0.801$\pm$0.039 0.831$\pm$0.042 0.847$\pm$0.004
Projection ours 0.770$\pm$0.011 0.798$\pm$0.008 0.804$\pm$0.004 0.757$\pm$0.004 0.793$\pm$0.003 0.738$\pm$0.001 0.816$\pm$0.016 0.823$\pm$0.020 0.853$\pm$0.002
Shen et al. 0.811$\pm$0.010 0.836$\pm$0.008 0.789$\pm$0.001 0.817$\pm$0.010 0.840$\pm$0.011 0.800$\pm$0.001 0.835$\pm$0.003 0.865$\pm$0.003 0.834$\pm$0.004
ACM Embedding ours 0.842$\pm$0.007 0.822$\pm$0.015 0.882$\pm$0.003 0.857$\pm$0.009 0.870$\pm$0.009 0.845$\pm$0.001 0.882$\pm$0.009 0.894$\pm$0.011 0.874$\pm$0.003